Circular No. 8952

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2008 L1 (LARSEN) R. S. McMillan, Lunar and Planetary Laboratory, University of Ari-

zona, reports that J. V. Scotti has recovered comet P/1997 V1 (cf. IAUC 6767) with the Spacewatch 1.8-m f/2.7 reflector at Kitt Peak (recovery observation tabulated below). On June 9.4 UT, Scotti described the coma as of diameter 9′′, and there was a faint tail extending 0′.15 in p.a. 243^{o}; on June 10.4, he gave the coma diameter as 10′′, the tail extending 0′.11 in p.a. 252^{o}. The indicated correction to the prediction on MPC 54170 (ephemeris on MPC 62064) is ΔT = −2.0 days.

2008 UT  α_{2000}  δ_{2000}  Mag.
June 9.44185  0^{h}25^{m}34.^{s}60  +13^{o}29′14′′.0  20.2

The available astrometry, the following orbital elements (along with revised elements for 1997), and an ephemeris appear on MPEC 2008-L45.

Epoch = 2008 Sept. 11.0 TT

T = 2008 Aug. 25.1744 TT  ω = 133.^{o}7286
e = 0.333226  Ω = 234.8163 }2000.0
q = 3.271999 AU  i = 12.1214

a = 4.907208 AU n^{o} = 0.0906677 P = 10.87 years

V5579 SAGITTARII R. J. Rudy, D. K. Lynch, R. W. Russell, K. Crawford, and B. Kane-

shiro, Aerospace Corporation; C. E. Woodward, University of Minnesota; M. Sitko, University of Cincinnati and Space Science Institute; and M. Skinner, Boeing LTS, report on spectroscopic observations (range 0.8–13.5 μm) of the nova V5579 Sgr from May 22 UT. Data from 0.8–5.2 μm were obtained using the SPEX instrument at the Infrared Telescope Facility, while measurements from 3–13.5 μm were acquired with the Advanced Electro-Optical Telescope using the Aerospace Corporation’s Broadband Array Spectrograph System. The nova has changed dramatically since the early observations from May 9 (IAUC 8948). The dust emission has increased and now dominates the continuum. The equivalent widths of the lines have weakened substantially, and the line profiles now show a pronounced absorption at their centers. The excitation of the emission lines remains low, although the He I line at 1.083 μm is now present. The dust has cooled significantly and now has a temperature of 1080 K.

2008 June 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green